OpenStudy (anonymous):

-35x^6y^4 OVER -7xy^5 how do i find this??

OpenStudy (anonymous):

\[\frac{ -35x ^{6}y ^{4} }{ -7xy ^{5} }=\frac{ 5x ^{5} }{ y }\] If you need me to explain, I will do so :D

OpenStudy (anonymous):

please do!

OpenStudy (anonymous):

Okay. We'll take it term by term and you'll understand it! -35/-7: This expression is simply division. And the negatives cancel, so 35/7 = 5. x^6/x: This expression has a higher exponent in the numerator. So we know x will remain in the numerator. Since we're dividing, subtract exponents. So 6-1 = 5, so x^5. y^4/y^5 This expression has a higher exponnt in the denominator. So we know y will remain in the denominator. Since we're dividing, subtract exponents. So 5-4 = 1, so y^1 or y. All together: 5x^5 / y

OpenStudy (anonymous):

ohh thanks again! i have another!

OpenStudy (anonymous):

haha bring it on :D

OpenStudy (anonymous):

When you have a simple equation like x+2> or equal to 5.... why do you switch it to less than or equal to?

OpenStudy (anonymous):

my bad i mean x+2< or = to 5. why do i switch it to greater than??